# Floer-theoretic filtration on Painlevé Hitchin systems

### v1

@@ Abstract.
We classify equivariant $\mathbb{C}^{*}$-actions on moduli spaces of Higgs bundles corresponding to the Painlevé equations. We deduce the Floer-theoretic filtrations on the cohomology of these spaces, introduced by Ritter and the second author in [RŽ23a]. Filtration is then compared with the “$P=W$” and the filtration obtained by multiplicities of the irreducible components of the nilpotent cone for all the 2-dimensional Higgs moduli.
@@ 1. Introduction
Hyperkähler moduli spaces of Higgs bundles on smooth compact curves, or with parabolic [^1] By parabolic Higgs bundles, we will always mean that the Higgs field has logarithmic poles compatible with a quasi-parabolic filtration. or irregular singularities, have recently seen numerous applications in various contexts in Mathematics and Physics. In the compact case, many such applications are made possible by the existence of a natural action of the multiplicative group $\mathbb{C}^{*}$ on the moduli spaces $\mathcal{M}$, by rescaling the Higgs field. Furthermore, this action is equivariant with respect to the Hitchin map
@@
In this article, we study equivariant holomorphic $\mathbb{C}^{*}$-actions on $2$-dimensional [^2] Dimensions, unless otherwise stated, are meant to be over the complex field. moduli spaces of Higgs bundles over curves. One set of examples correspond to the Painlevé equations $PI,\ldots,PVI$ (see Section 3). An extensive literature is devoted to these spaces from various perspectives. Here, we give a necessary and sufficient condition for the existence of a $\mathbb{C}^{*}$-action on these spaces $\mathcal{M}^{PX}$ (see Proposition 3.10), and spell out explicitly the construction of such an action whenever it exists. Furthermore, we compute its weights on the Hitchin base $\mathcal{B}^{PX}$ and on the symplectic $2$-form $\Omega_{I}^{PX}$ on $\mathcal{M}^{PX}.$ Another source of examples are moduli spaces $\mathcal{M}_{\Gamma}$ of parabolic Higgs bundles corresponding to the affine root systems $\widetilde{D}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8}$, see for instance [Boa18, Section 4.1][Osh13, Proposition 8.1][Gro14][Zha17]. A degenerate case in this family corresponding to $\widetilde{A}_{0}$ is the cotangent bundle of an elliptic curve. We carry out the same study for these spaces too. The topology of the case $\widetilde{E}_{6}$ has been studied in detail in [Mih23].
As an example of the general theory developed in [RŽ23a], the constructed $\mathbb{C}^{*}$-actions endow the cohomology spaces with the structure of a filtration $\mathscr{F}_{\bullet}$ by using Floer-theoretic ideas. We review key features of this construction in Section 2. Given a label $X$ in the Painlevé classification, we introduce a two-variable generating Puiseux polynomial $P^{PX}$ that encodes the non-trivial weights of $\mathscr{F}$ on the cohomology of the corresponding moduli space $\mathcal{M}^{PX}$ along with their dimensions:
@@
where $\mathbb{B}$ is any base field. Our first result may be summarized as:
@@
Similar actions from a little bit different perspective are constructed in [FPYY18, Appendix C]. The main difference is that our approach comes with a complete geometric understanding of the underlying moduli spaces, allowing us to work out the full Morse theory picture. On the other hand, [FPYY18, Appendix C] works for higher dimensional moduli spaces of irregular Higgs bundles of rank $2$ in genus 00 too. We hope that the combination of these two approaches can bring new insights into the topology of such moduli spaces. A potential application of our result is towards Mirror Symmetry. Indeed, using the terminology of [KW07, Section 5.5], the flowlines of our $\mathbb{C}^{*}$-actions are branes of type $(B,A,A)$, that are expected to have duals of type $(B,B,B)$.
The computation of filtration $\mathscr{F}$ was done in [RŽ23b] for the (aforementioned) 2-dimensional parabolic Higgs moduli $\mathcal{M}_{\Gamma}$, where $\Gamma=\{0,\mathbb{Z}/2,\mathbb{Z}/3,\mathbb{Z}/4,\mathbb{Z}/6\}$. To groups $\Gamma$ one associates affine Dynkin graphs $Q_{\Gamma}:=\widetilde{A}_{0},\widetilde{D}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8}$, respectively. The intersection between these and Painlevé spaces is precisely $\mathcal{M}_{\mathbb{Z}/2}=\mathcal{M}^{PVI}$. It has been shown that:
@@
[RŽ23b] $\mathscr{F}(H^{2}(\mathcal{M}_{\Gamma}))$ refines the $P=W$ filtration. Moreover, its ranks correspond to the labels of the imaginary root of the graph $Q_{\Gamma}.$
@@
The labellings of the imaginary root mentioned in 1.2 can be described in terms of the multiplicities of the irreducible components of the central fiber $h_{\Gamma}^{-1}(0)$ of the Hitchin map
@@
with its inherited scheme structure. That motivates us to define multiplicity filtration, for general Higgs moduli $\mathcal{M}$. Given the central fiber of its Hitchin map with its irreducible components $E_{i}$ and their multiplicities $m_{i},$
@@
Interestingly, the multiplicities $m_{i}$ are computed for irreducible components $E_{i}$ of a special type (called very stable Higgs bundles) in recent work by Hausel–Hitchin [HH22], where they show that these numbers show up in the mirror symmetry for Higgs moduli spaces.
@@
Given a $2$-dimensional Higgs moduli space $\mathcal{M}$ which has a $I$-holomorphic $\mathbb{C}^{*}$-action and such that the Hitchin map has generically smooth fibers, its multiplicity and Floer-theoretic filtration coincide rank-wise. For moduli spaces of parabolic Higgs bundles, they both refine the $P=W$ filtration, whereas for Painlevé spaces, $P=W$ refines the former two filtrations. Thus, in the intersection, i.e. for Painlevé VI space, these three filtrations coincide.
Thus, the last theorem compares the three filtrations for all 2-dimensional Higgs moduli for which filtrations $\mathscr{F}$ and $\mathscr{M}$ make sense. An interesting avenue for future research is to understand how this generalizes in higher dimensions. Due to [Gro14], we know that the Hilbert schemes of parabolic 2-dimensional moduli, $\mathrm{Hilb}^{n}(\mathcal{M}_{\Gamma})$, are parabolic Higgs moduli spaces as well. It would be interesting to understand the relation between the $P=W$, Floer-theoretic, and multiplicity filtration for these spaces.
Acknowledgements. We thank Alexandre Minets for helpful conversations, in particular for pointing out the coincidence between the multiplicities of the components of central fiber and corresponding imaginary root, for 2-dimensional parabolic Higgs moduli. This work grew out of discussions at the workshop ”Birational Geometry and Quantum Invariants” held at the Simons Center for Geometry and Physics in November 2023. The first author would like to thank the organizers of the meeting for the invitation, the Center for its hospitality, and grants KKP 144148 and K146401 of the agency NKFIH (Hungary) for support. The second author wishes to thank the Simons Center for its hospitality.
@@ 2. Filtration from the Floer theory of a $\mathbb{C}^{*}$-action
In this section, we recall briefly the relevant results from [RŽ23a, RŽ23b], to which we refer an interested reader to look for details. In that work, the authors consider connected symplectic manifolds $(Y,\omega)$ admitting a contracting $\mathbb{C}^{*}$-action $\varphi$, which means that given an arbitrary point $y\in Y$, convergence point
@@
always exists, and the set of these points, equal to the set of fixed points $F:=Y^{\mathbb{C}^{*}}$, is compact. In addition, the $\mathbb{C}^{*}$-action has to be $I$-pseudoholomorphic with respect to an almost complex structure $I$ compatible with symplectic form $\omega.$ Furthermore, the $S^{1}$-part has to be Hamiltonian, meaning that there is a moment map
$H:Y\rightarrow\mathbb{R}.$
Altogether, this is what is called a symplectic $\mathbb{C}^{*}$-manifold, which we assume for $Y$ from now on. Such manifolds have an important compact subset called the core
@@
to an affine space with a linear $\mathbb{C}^{*}$-action with only positive weights, then $\mathrm{Core}(Y)=\Psi^{-1}(0).$
Consider a directed system of Hamiltonian Floer cohomologies $\{HF^{*}(H_{\lambda})\}_{\lambda\in\mathbb{R}\setminus\mathcal{P}}$ with Hamiltonians $H_{\lambda}$ that are, outside a compact set, linear with slope $\lambda$ with respect to $H$, and $\lambda$ does not belong to the (discrete) set of periods $\mathcal{P}$ of the $S^{1}$-action $\varphi$. The authors in [RŽ23a] define the filtration by ideals on quantum cohomology ring $QH^{*}(Y)$:
@@
where $c_{\lambda}^{*}$ is a Floer-theoretic type of map (called a continuation map). For $p\in\mathcal{P},$ simply define $\mathscr{F}_{p}^{\varphi}:=\cap_{p<\lambda}\mathscr{F}_{\lambda}^{\varphi}.$ Quantum cohomology is $QH^{*}(Y)=H^{*}(Y,\mathbb{K})$ as a vector space over the Novikov field
(6) $\mathbb{K}=\{\sum n_{j}T^{a_{j}}\mid a_{j}\in\mathbb{R},a_{j}\to\infty,n_{j}\in\mathbb{B}\},$
where $T$ is a formal variable in grading zero, and $\mathbb{B}$ is any choice of base field. Therefore, by “letting $T\rightarrow 0$”, one defines the specialised $\varphi$-filtration on ordinary cohomology ring $H^{*}(Y,\mathbb{B})$ by cup-ideals
@@
Filtration (7) is thus more feasible than (5) to be comparable to other known filtrations, as we will see in the examples. However, in practice, we will often blur the difference between them in practice as
@@
For the spaces considered in this paper, these are easily-computable, due to333for proof see e.g. [RŽ23a, Lemma 8.7]
@@
If $Y$ has an $I$-holomorphic symplectic structure $\Omega_{I}$ with positive $\mathbb{C}^{*}$-weight, $t\cdot\Omega_{I}=t^{s}\Omega_{I}$ for some integer $s>0,$ then there is a non-degenerate pairing
@@
The $S^{1}$-moment map $H$ is a Morse–Bott function, and one has a decomposition444where the shift notation is usual – given a graded module $A^{*},$ the graded module $A^{*}[k]$ is defined by $A[k]^{i}:=A^{i+k}.$
@@
Let $\lambda_{\mathrm{min}}:=1/(\textrm{maximal absolute weight of }F_{\mathrm{min}})$. If $H^{*}(Y)$ lies in even degrees, and $\lambda<\lambda_{\mathrm{min}}$, then $c_{\lambda}^{*}:H^{*}(F_{\mathrm{min}})\to HF^{*}(H_{\lambda})$ is injective, so $\mathscr{F}_{\lambda}\cap H^{*}(F_{\mathrm{min}})=\{0\}$, and in (10)
$\mathscr{F}_{\mathbb{B},\lambda}\subset\oplus_{\alpha\neq\mathrm{min}}H^{*}(F_{\alpha};\mathbb{B})[-\mu_{\alpha}].$
Moreover, there is a statement describing when the unit $1\in H^{0}(F_{\mathrm{min}})=H^{0}(Y)$ enters the filtration, in the setup of 2.3:
@@
On the other hand, there are lower bounds for ranks of the filtration, which are readily-computable, using only the Betti numbers $b_{k}(F_{\alpha})$ of the fixed loci, and their weight decompositions (9). Introduce Floer-theoretic indices
(11) $\mu_{\lambda}(F_{\alpha})=\dim_{\mathbb{C}}Y-\dim_{\mathbb{C}}F_{\alpha}-\sum_{k}\dim_{\mathbb{C}}(H_{k})W(\lambda k),$
@@
The statements mentioned so far completely determine the filtration in the case of Painlevé spaces. However, in order to determine the filtration in general, one has to use the Morse–Bott–Floer spectral sequences defined in [RŽ23b]. Let us briefly describe this method. Given an integer $m>1$, observe the fixed locus $Y_{m}:=Y^{\mathbb{Z}/m}$ of the cyclic subgroup $\mathbb{Z}/m\leq\mathbb{C}^{*}$. It is an $I$-pseudoholomorphic symplectic submanifold that breaks into a finite number of connected components, out of which we label by $\{Y_{m,\beta}\}_{\beta}$ the non-compact ones, or equivalently those that go out of the $\mathrm{Core}(Y).$555recall the definition of core in (3) We call them outer torsion submanifolds. Denote their intersections with energy hypersurfaces by
@@
(14) $E_{1}^{pq}(\varphi)=\begin{cases}H^{q}(Y)&\text{if }p=0\textrm{ and }\\ \bigoplus_{\beta}H_{p,\beta}^{*}[-\mu(B_{T_{p},\beta})]&\text{if }p\leq 0\\ 0&otherwise.\end{cases}$
This spectral sequence determines the filtration: a class $x\in H^{*}(Y)=QH^{*}(Y)$ lies in $\mathscr{F}^{\varphi}_{\lambda}$ iff the columns of the spectral sequence having $T_{p}\leq\lambda$ annihilate $x\in E_{1}^{0,*}=H^{*}(Y)$ via some differential $\operatorname{d}_{r}$ of the spectral sequence $E_{r}^{pq}(\varphi)$.
@@
where $w_{i}$ are the weights of the convergence point666recall (2) $y_{0}$ of an arbitrary $y\in B_{T_{p},\beta}.$ In general, there is another Floer-theoretic local energy spectral sequence
$\oplus_{\beta}H^{*}(B_{T_{p},\beta})[-\mu(B_{T_{p},\beta})]\Rightarrow\oplus_{\beta}H_{p,\beta}^{*}[-\mu(B_{T_{p},\beta})]$
which collapses in low dimensions $\dim_{\mathbb{C}}(Y)\leq 2,$ in particular for all spaces considered in this paper, so computing (14) amounts to computing ordinary cohomologies $H^{*}(B_{T_{p},\beta}),$ and their shifts $\mu(B_{T_{p},\beta})$. In practice one computes these by knowing the topology of fixed loci $F_{\alpha}$ and their weight decompositions (9), as each $Y_{m,\beta}$ is obtained via the $\mathbb{C}^{*}$-flow from (one or more) fixed $F_{\alpha},$ where it converges (when $t\rightarrow 0$) to the ($\mathbb{Z}/m$-fixed) part
@@
3. Construction of actions and filtration computation

### v2

@@ Abstract.
We classify equivariant $\mathbb{C}^{*}$-actions on moduli spaces of Higgs bundles corresponding to the Painlevé equations. Using this, we compute the Floer-theoretic filtrations on the cohomology of these spaces, introduced by Ritter and the second author in [RŽ23a]. We compare it with the “$P=W$” and the filtration obtained by multiplicities of the irreducible components of the nilpotent cone, ultimately deducing that the Floer-theoretic filtration coincides with the multiplicity filtration, for all 2-dimensional Higgs moduli.
@@ 1. Introduction
Hyperkähler moduli spaces of Higgs bundles on smooth compact curves, possibly with parabolic [^1] By parabolic Higgs bundles, we will always mean that the Higgs field has logarithmic poles compatible with a quasi-parabolic filtration. or irregular singularities, have recently seen numerous applications in various contexts in Mathematics and Physics. In the compact (i.e., non-punctured) case, many such applications are made possible by the existence of a natural action of the multiplicative group $\mathbb{C}^{*}$ on the moduli spaces $\mathcal{M}$, by rescaling the Higgs field. Furthermore, this action is equivariant with respect to the Hitchin map, defined as the characteristic polynomial of the Higgs field
@@
In this article, we study equivariant holomorphic $\mathbb{C}^{*}$-actions on $2$-dimensional [^2] Dimensions, unless otherwise stated, are meant to be over the complex field. moduli spaces of Higgs bundles over curves. One set of examples corresponds to the Painlevé equations $PI,\ldots,PVI$ (see Section 3). Extensive literature is devoted to these spaces from various perspectives. Here, we give a necessary and sufficient condition for the existence of a $\mathbb{C}^{*}$-action on these spaces $\mathcal{M}^{PX}$, and spell out explicitly the construction of such an action whenever it exists. Furthermore, we compute its weights on the Hitchin base $\mathcal{B}^{PX}$ and on the symplectic $2$-form $\Omega_{I}^{PX}$ on $\mathcal{M}^{PX}.$ Another source of examples are moduli spaces $\mathcal{M}_{\Gamma}$ of parabolic Higgs bundles corresponding to the affine root systems $\widetilde{A}_{0},\widetilde{D}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8}$, see for instance [Boa18, Section 4.1][Osh13, Proposition 8.1][Gro14][Zha17]. A degenerate case in this family corresponding to $\widetilde{A}_{0}$ is the cotangent bundle of an elliptic curve. We carry out the same study for these spaces, too. The topology of the case $\widetilde{E}_{6}$ has been studied in detail in [Mih23].
As an example of the general theory developed in [RŽ23a] by using Floer-theoretic techniques, the constructed $\mathbb{C}^{*}$-actions endow the singular cohomology of these spaces with the structure of a filtration $\mathscr{F}_{\bullet}$. We review key features of this construction in Section 2. Given a label $X$ in the Painlevé classification, we introduce a two-variable generating Puiseux polynomial $P^{PX}$ that encodes the non-trivial weights of $\mathscr{F}$ on the cohomology of the corresponding moduli space $\mathcal{M}^{PX}$ along with their dimensions:
@@
where $\mathbb{B}$ is any base field of characteristic zero. Our first result may be summarized as:
@@
Similar actions from a little bit different perspective are constructed in [FPYY18, Appendix C]. The main difference is that our approach comes with a complete geometric understanding of the underlying moduli spaces, allowing us to work out the full Morse theory picture. On the other hand, [FPYY18, Appendix C] works for higher-dimensional moduli spaces of irregular Higgs bundles of rank $2$ in genus 00 too. We hope that the combination of these two approaches can bring new insights into the topology of such moduli spaces. A potential application of our result is towards Mirror Symmetry. Indeed, using the terminology of [KW07, Section 5.5], the flowlines of our $\mathbb{C}^{*}$-actions are branes of type $(B,A,A)$, that are expected to have duals of type $(B,B,B)$.
The computation of filtration $\mathscr{F}$ was done in [RŽ23b] for the (aforementioned) 2-dimensional parabolic Higgs moduli $\mathcal{M}_{\Gamma}$, where $\Gamma=\{0,\mathbb{Z}/2,\mathbb{Z}/3,\mathbb{Z}/4,\mathbb{Z}/6\}$. To groups $\Gamma$, one associates affine Dynkin graphs $Q_{\Gamma}:=\widetilde{A}_{0},\widetilde{D}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8}$, respectively. The intersection between these and Painlevé spaces is precisely $\mathcal{M}_{\mathbb{Z}/2}=\mathcal{M}^{PVI}$. It has been shown that:
@@
[RŽ23b] The Floer-theoretic filtration $\mathscr{F}(H^{2}(\mathcal{M}_{\Gamma}))$ refines the $P=W$ filtration. Moreover, its ranks correspond to the labels of the imaginary root of the graph $Q_{\Gamma}.$
@@
The labelling of the imaginary root mentioned in 1.2 can be described in terms of the multiplicities of the irreducible components of the central fiber $h_{\Gamma}^{-1}(0)$ of the Hitchin map
@@
with its inherited scheme structure. That motivates us to define multiplicity filtration, for a general Higgs moduli space $\mathcal{M}$. Given the central fiber of its Hitchin map with its irreducible components $E_{i}$ and their multiplicities $m_{i},$
@@
Interestingly, the multiplicities $m_{i}$ are discussed for the irreducible components $E_{i}$ of a special type (called very stable Higgs bundles) in recent work by Hausel–Hitchin [HH22], where they show that these numbers show up in the (conjectured) mirror symmetry for Higgs moduli spaces.
@@
Given a $2$-dimensional Higgs moduli space $\mathcal{M}$ which has a holomorphic $\mathbb{C}^{*}$-action and such that the Hitchin map has generically smooth fibers, its multiplicity and Floer-theoretic filtration coincide rank-wise
$\mathrm{rk}\,\mathscr{M}_{\lambda}=\mathrm{rk}\,\mathscr{F}_{\lambda}.$
When $\mathcal{M}$ is a parabolic Higgs moduli, they both refine the $P=W$ filtration, whereas for Painlevé spaces, $P=W$ refines the former two filtrations. Thus, in the intersection, i.e. for Painlevé VI space, these three filtrations coincide.
Thus, the last theorem compares the three filtrations for all 2-dimensional Higgs moduli for which filtrations $\mathscr{F}$ and $\mathscr{M}$ make sense. An interesting avenue for future research is to understand how this generalises in higher dimensions. Due to [Gro14], we know that the Hilbert schemes of parabolic 2-dimensional moduli, $\mathrm{Hilb}^{n}(\mathcal{M}_{\Gamma})$, are parabolic Higgs moduli spaces as well. It would be interesting to understand the relation between the $P=W$, Floer-theoretic, and multiplicity filtration for these spaces.
Acknowledgments. We thank Alexandre Minets for helpful conversations, in particular for pointing out the coincidence between the multiplicities of the components of the central fiber and the corresponding imaginary root, for 2-dimensional parabolic Higgs moduli. This work grew out of discussions at the workshop ”Birational Geometry and Quantum Invariants” held at the Simons Center for Geometry and Physics in November 2023. The first author would like to thank the organizers of the meeting for the invitation, the Center for its hospitality, and grants KKP 144148 and K146401 of the agency NKFIH (Hungary) for support. The second author wishes to thank the Simons Center for its hospitality.
@@ 2. Filtration from the Floer theory of a $\mathbb{C}^{*}$-action
In this section, we recall the relevant results from the construction of the Floer-theoretic filtration by Ritter–Živanović, which we are going to use in order to compute this filtration in the next section. To keep the current paper brief, we do not include many details regarding the Floer theory, but rather we refer the interested reader to their papers [RŽ23a, RŽ23b].
In that work, the authors consider connected symplectic manifolds $(Y,\omega)$ admitting a contracting $\mathbb{C}^{*}$-action $\varphi$, which means that given an arbitrary point $y\in Y$, convergence point
@@
always exists, and the set of these points, equal to the set of fixed points $F:=Y^{\mathbb{C}^{*}}$, is compact. The $\mathbb{C}^{*}$-action is $I$-pseudoholomorphic with respect to an almost complex structure $I$ compatible [^3] Meaning that $\omega(\cdot,I\cdot)$ gives a Riemannian metric. with symplectic form $\omega.$ Furthermore, the $S^{1}$-part has to be Hamiltonian, meaning that there is a moment map, i.e. a function
$H:Y\rightarrow\mathbb{R}$
satisfying $\omega(\cdot,X_{S^{1}})=dH,$ where $X_{S^{1}}$ is the vector-field of the $S^{1}$-action. Altogether, this is what is called a symplectic $\mathbb{C}^{*}$-manifold, which we assume for $Y$ from now on. Such manifolds have an important compact subset called the core
@@
to an affine space $\mathcal{B}$ with a linear $\mathbb{C}^{*}$-action with only positive weights, then $\mathrm{Core}(Y)=\Psi^{-1}(0).$
Consider a directed system of Hamiltonian Floer cohomologies $\{HF^{*}(H_{\lambda})\}_{\lambda\in\mathbb{R}\setminus\mathcal{P}}$ with Hamiltonians $H_{\lambda}$ that are, outside a compact set, linear with slope $\lambda$ with respect to $H$, and $\lambda$ does not belong to the (discrete) set of periods $\mathcal{P}$ of the $S^{1}$-action $\varphi$.
Briefly, $HF^{*}(H_{\lambda})$ is the homology of the chain complex generated by 1-periodic orbits of the Hamiltonian vector-field $X_{H_{\lambda}}$ (defined as the dual of the differential, $\omega(\cdot,X_{H_{\lambda}})=dH_{\lambda}$) and the differential counts pseudo-holomorphic curves perturbed by the same vector field, so satisfying the equation $\partial_{s}u+I(\partial_{t}u-X_{H_{\lambda}})=0.$ The cohomological gradings are by certain Robbin–Salamon indices, which in general are hard to compute. But in our setup, since the $S^{1}$-action is pseudoholomorphic, it amounts to computing winding numbers of paths of unitary matrices, which is computable, as we will see (Equations (11), (15)).
The authors in [RŽ23a] define the filtration by ideals on quantum cohomology ring $QH^{*}(Y)$:
@@
where $c_{\lambda}^{*}$ is a Floer-theoretic type of map (called a continuation map). For $p\in\mathcal{P},$ simply define $\mathscr{F}_{p}^{\varphi}:=\cap_{p<\lambda}\mathscr{F}_{\lambda}^{\varphi}.$ Quantum cohomology as a $\mathbb{K}$-vector space is $QH^{*}(Y)=H^{*}(Y,\mathbb{K})$, with
(6) $\mathbb{K}:=\{\sum n_{j}T^{a_{j}}\mid a_{j}\in\mathbb{R},a_{j}\to\infty,n_{j}\in\mathbb{B}\},$
is the Novikov field, with formal variable $T$ in grading zero, and $\mathbb{B}$ is any choice of base field of characteristic zero. [^4] This condition is not necessary for the theory, but it is assumed in the computations, to avoid dealing with torsion coefficients. Therefore, by “letting $T\rightarrow 0$”, one defines the specialized $\varphi$-filtration on ordinary cohomology ring $H^{*}(Y):=H^{*}(Y,\mathbb{B})$ by cup-ideals
@@
Filtration (7) is thus more feasible than (5) to be comparable to other known filtrations, as we will see in the examples. However, in practice, we will often blur the difference between them in practice as (by linear algebra) we have
@@
For the spaces considered in this paper, these decompositions are easily-computable, due to:555for proof see e.g. [RŽ23a, Lemma 8.7]
@@
Fix a component $F_{\alpha}$. If $Y$ has an $I$-holomorphic symplectic structure $\Omega_{I}$ with positive $\mathbb{C}^{*}$-weight, $t\cdot\Omega_{I}=t^{s}\Omega_{I}$ for some integer $s>0,$ then there is a non-degenerate pairing
@@
The $S^{1}$-moment map $H$ is a Morse–Bott function, and one has a decomposition666where the shift notation is usual – given a graded module $A^{*},$ the graded module $A^{*}[k]$ is defined by $A[k]^{i}:=A^{i+k}.$
@@
Let $\lambda_{\mathrm{min}}:=1/(\textrm{maximal absolute weight of }F_{\mathrm{min}})$. If $H^{odd}(Y)=0$, and $\lambda<\lambda_{\mathrm{min}}$, then $c_{\lambda}^{*}:H^{*}(F_{\mathrm{min}};\mathbb{K})\to HF^{*}(H_{\lambda})$ is injective, so $\mathscr{F}_{\lambda}\cap H^{*}(F_{\mathrm{min}};\mathbb{K})=\{0\}$, and in (10)
$\mathscr{F}_{\mathbb{B},\lambda}\subset\oplus_{\alpha\neq\mathrm{min}}H^{*}(F_{\alpha})[-\mu_{\alpha}].$
Moreover, there is a statement describing when the unit $1\in H^{0}(F_{\mathrm{min}})=H^{0}(Y)$ enters the filtration (i.e. smallest $\lambda$ such that $1\in\mathscr{F}_{\lambda}$), in the setup of 2.3:
@@
On the other hand, there are lower bounds for ranks of the filtration, which are readily computable, using only the Betti numbers $b_{k}(F_{\alpha})$ of the fixed loci, and their weight decompositions (9). Introduce Floer-theoretic indices
(11) $\mu_{\lambda}(F_{\alpha}):=\dim_{\mathbb{C}}Y-\dim_{\mathbb{C}}F_{\alpha}-\sum_{k}\dim_{\mathbb{C}}(H_{k})W(\lambda k),$
@@
The statements mentioned so far completely determine the filtration in the case of Painlevé spaces. However, in order to determine the filtration in general, one has to use the Morse–Bott–Floer spectral sequences defined in [RŽ23b]. Let us briefly describe this method. Given an integer $m>1$, observe the fixed locus $Y_{m}:=Y^{\mathbb{Z}/m}$ of the cyclic subgroup $\mathbb{Z}/m\leq\mathbb{C}^{*}$. It is an $I$-pseudoholomorphic symplectic submanifold that breaks into a finite number of connected components, out of which we label by $\{Y_{m,\beta}\}_{\beta}$ the non-compact ones, or equivalently those that go out of the $\mathrm{Core}(Y).$777recall the definition of core in (3) We call them outer torsion submanifolds, as they have torsion isotropy subgroups of the $\mathbb{C}^{*}$-action. Denote their intersections with energy hypersurfaces by
@@
(14) $E_{1}^{pq}(\varphi)=\begin{cases}H^{q}(Y;\mathbb{K})&\text{if }p=0\textrm{ and}\\ \bigoplus_{\beta}H_{p,\beta}^{*}[-\mu(B_{T_{p},\beta})]&\text{if }p\leq 0\\ 0&otherwise.\end{cases}$
This spectral sequence determines the filtration: a class $x\in H^{*}(Y;\mathbb{K})=QH^{*}(Y)$ lies in $\mathscr{F}^{\varphi}_{\lambda}$ iff the columns of the spectral sequence having $T_{p}\leq\lambda$ annihilate $x\in E_{1}^{0,*}=H^{*}(Y;\mathbb{K})$ via some differential $\operatorname{d}_{r}$ of the spectral sequence $E_{r}^{pq}(\varphi)$.
@@
where $w_{i}$ are the weights of the convergence point888recall (2) $y_{0}$ of an arbitrary $y\in B_{T_{p},\beta}.$ In general, there is another Floer-theoretic local energy spectral sequence
$\oplus_{\beta}H^{*}(B_{T_{p},\beta};\mathbb{K})[-\mu(B_{T_{p},\beta})]\Rightarrow\oplus_{\beta}H_{p,\beta}^{*}[-\mu(B_{T_{p},\beta})]$
which collapses in low dimensions $\dim_{\mathbb{C}}(Y)\leq 2,$ in particular for all spaces considered in this paper, so computing (14) amounts to computing ordinary cohomologies $H^{*}(B_{T_{p},\beta};\mathbb{K}),$ and their shifts $\mu(B_{T_{p},\beta})$. In practice, one computes these by knowing the topology of fixed loci $F_{\alpha}$ and their weight decompositions (9), as each $Y_{m,\beta}$ is obtained via the $\mathbb{C}^{*}$-flow from (one or more) fixed $F_{\alpha},$ where it converges (when $t\rightarrow 0$) to the ($\mathbb{Z}/m$-fixed) part
@@
3. Construction of $\mathbb{C}^{*}$-actions and filtration computation